# Revisiting the Thickness of the Air-Water Interface from Two Extremes of Interface Hydrogen Bond Dynamics

### v2

@@ Abstract
The air-water interface plays a crucial role in many aspects of science, because of its unique properties, such as a two-dimensional hydrogen bond (HB) network and completely different HB dynamics compared to bulk water. However, accurately determining the boundary of interfacial and bulk water, that is, the thickness of the air-water interface, still challenges experimentalists. Various simulation-based methods have been developed to estimate the thickness, converging on a range of approximately 3 to 10 Å. In this study, we introduce a novel approach, grounded in density functional theory-based molecular dynamics (DFTMD) simulations, to measure the air-water interface thickness, offering a different perspective based on prior research. To capture realistic HB dynamics in the air-water interface, two extreme scenarios of the interface HB dynamics are obtained: one underestimates the interface HB dynamics, while the other overestimates it. Surprisingly, our results suggest that the interface HB dynamics in both scenarios converge as the thickness of the air-water interface increases to 4.0 Å. This convergence point, indicative of the realistic interface thickness, is confirmed by our calculation of anisotropic decay of OH stretch at the air-water interface and supported by the results from recent ab initio molecular dynamics simulations.
@@ I Introduction
Advances in the study of hydrogen bond (HB) dynamics at the air-water interface have been significant. Liu et al. Liu, Harder, and Berne (2005) used molecular dynamics (MD) simulations to demonstrate faster HB breaking and forming at the interface than bulk water, attributed to quicker translational diffusion. From sum frequency generation (SFG) vibrational spectroscopy, Gan et al. Gan et al. (2006a) found that at the air-water interface, singly hydrogen (H)-bonded water molecules align almost parallel to the interface with limited orientational variation, while doubly H-bonded donor molecules orient their dipole vectors away from the liquid phase, highlighting diverse behaviors among interfacial water molecules. Almost concurrently, through time-resolved SFG vibrational spectroscopy, McGuire and ShenMcguire and Shen (2006) observed ultrafast vibrational dynamics at the interface, noting that the relaxation behaviors of interfacially bonded OH stretch modes on sub-picosecond timescales were akin to those in bulk water, encompassing spectral diffusion, vibrational relaxation, and thermalization. Pioneering work by Tahara’s group Singh et al. (2013); Inoue et al. (2016), which presented the first two-dimensional heterodyne-detected vibrational SFG (2D HD-VSFG) spectra of the OH stretch region at the interface, highlighted diverse behaviors among interfacial HB OH groups. Subsequent studies, including those by Jeon et al. Jeon et al. (2017) and Ojha and Kühne Ojha and Kühne (2021), have used MD and ab initio MD (AIMD) simulations to explore the structure and dynamics of interfacial water, uncovering weaker H-bonds and faster vibrational spectral dynamics of free OH groups compared to H-bonded OH groups at the interface. These collective insights enhance our understanding of the vibrational energy relaxation, HB dynamics, and interactions of water molecules at the air-water interface. Building upon this knowledge of interfacial behavior, significant efforts have also been directed towards quantifying the physical characteristics of the interface, including its thickness, which plays a crucial role in understanding its molecular interactions and behavior.
@@
The air-water interface thickness has been measured via ellipsometry Raman and Ramdas (1927); McBain, Bacon, and Bruce (1939); Kinosita and Yokota (1965), relative permittivity measurements Fumagalli et al. (2018), X-ray reflectivity Braslau et al. (1985, 1988), SFG spectroscopy Superfine, Huang, and Shen (1991); Du et al. (1993); Ostroverkhov, Waychunas, and Shen (2005); Groenzin et al. (2007); Shen (2012); Ishiyama, Takahashi, and Morita (2012); Shen (2013); Singh et al. (2013); Ohto et al. (2015); Tang et al. (2020), classical MD simulationsLiu, Harder, and Berne (2005); Kuo et al. (2006); Wick et al. (2007); Morita and Garrett (2008); Sedlmeier, Horinek, and Netz (2009); Fan et al. (2009); Ishiyama and Morita (2009); Willard and Chandler (2010); Vila Verde, Bolhuis, and Campen (2012); Ishiyama, Morita, and Tahara (2015); Nagata et al. (2015) and the AIMD simulationsKühne et al. (2011); Nagata et al. (2012); Kessler et al. (2015); Jeon et al. (2017); Ohto et al. (2019); Tang et al. (2020) to mention just a few. There is a consensus that the thickness of the air-water interface is about 3 to 10 Å Jungwirth and Tobias (2006); Sedlmeier, Horinek, and Netz (2009); Ishiyama and Morita (2009); Kuo et al. (2006); Wick et al. (2007); Fan et al. (2009); Kühne et al. (2011); Baer et al. (2014); Kessler et al. (2015); Nguyen, Nakahara, and Phan (2020); Peng et al. (2021). Nonetheless, accurately determining the thickness remains experimentally challenging. The MD and Monte Carlo (MC) simulations of the air-water interface yield molecular-level information not readily available in experiments. These simulations, which utilize various intermolecular potential functions, have played a crucial role in estimating the thickness Townsend, Gryko, and Rice (1985); Wilson, Pohorille, and Pratt (1987); Matsumoto and Kataoka (1988); Townsend and Rice (1991); Lie et al. (1993); Alejandre, Tildesley, and Chapela (1995); Taylor, Dang, and Garrett (1996); Dang and Chang (1997b); Tang et al. (2020). Additionally, density functional theory-based MD (DFTMD) simulations Car and Parrinello (1985); Marx and Hutter (2000); Car (2002); Kuo and Mundy (2004); Kühne et al. (2007, 2011); Khaliullin and Kühne (2013) also offer a predictive platform for understanding density profiles and determining the thickness of the air-liquid interfaces Blas et al. (2001); Kühne et al. (2011); Kessler et al. (2015); Pezzotti, Galimberti, and Gaigeot (2017).
@@
Inspired by the above experimental and simulation results, and with the motivation of capturing realistic HB dynamics at interfaces, we have designed an approach based on two extreme scenarios of interface HB dynamics. In the first scenario, for the set of molecules located in the interface layer at a given sampling time, we use the Luzar-Chandler (LC) HB population operator Luzar and Chandler (1996a) to obtain the HB dynamics of these interface molecules. In the second scenario, taking inspiration from Luzar and Chandler’s HB population and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017), we have developed an interface HB (IHB) population operator. This operator aims to provide a refined understanding of HB dynamics specifically at the interface.
@@
Due to the thermal motion of water molecules, two scenarios may occur. In the first scenario, the water molecules under observation might transition into the bulk phase. In the second scenario, if a water molecule resides in the interface region and its H-bonded partner moves outside the interface area, then such a pair of molecules will no longer be of concern. Based on the study of interfacial water dynamics by Liu et al. Liu, Harder, and Berne (2005), Gan et al. Gan et al. (2006a), Singh et al. Singh et al. (2013) and Jeon et al. Jeon et al. (2017), as well as the investigation into the time-dependent spectral evolution of H-bonded and free water molecules by Ojha and Kühne Ojha and Kühne (2021), our approach indicates that the HB dynamics derived from the first scenario will be slower the genuine interfacial HB dynamics. Conversely, the one obtained from the second scenario will exhibit a faster dynamics than the genuine one.
Building upon the foundation laid by methods reliant on the density criterion Townsend and Rice (1991); Beaglehole and Wilson (1993); Lie et al. (1993); Taylor, Dang, and Garrett (1996); Sokhan and Tildesley (1997); Alejandre, Tildesley, and Chapela (1995); Sokhan and Tildesley (1997); Morita and Hynes (2000); Paul and Chandra (2004); Morita (2006); Zhao et al. (2018), our approach introduces an alternative way of determining interface thickness through the analysis of the convergence of interfacial HB dynamics properties. This approach effectively bypasses the necessity of accounting for liquid density. As such, it offers another perspective for measuring the thickness of air-water interfaces. Furthermore, the principles underlying our approach hold potential for application to a broader range of systems, such as solution interfaces and ion shells, offering a flexible tool for interface studies.
@@ II Methods
Figure 1: A slab of water with the instantaneous surface $\mathbf{S}$ represented as a blue mesh on the upper and lower phase boundary. The grey surface, which represents an imaginary interface $\mathbf{S}^{\prime}$, is obtained by translating the surface to the inside of the system along the normal ($z$-axis) by distance $d$. This distance $d$ is then utilized to measure the thickness of the air-water interface.
For the slab in the cuboid simulation box, an imaginary surface ${\mathbf{S}}^{\prime}(t)$ is obtained by translating the surface ${\mathbf{S}}(t)$ along the system’s normal (into bulk) to a distance $d$. The region between the two surfaces ${\mathbf{S}}(t)$ and ${\mathbf{S}}^{\prime}(t)$ is defined as the air-water interface. Below we will combine two extreme scenarios to investigate the HB dynamics at the instantaneous air-water interface.
II.1 Scenario 1: HB dynamics based on the Luzar-Chandler HB population $h[{\bf r}(t)]$
As the first scenario we use the Luzar-Chandler HB population operator Luzar and Chandler (1996a) and employ a technique that samples water molecules right at the instantaneous interface for certain sampling time points. During the $t_{\rm traj}$ period, most water molecules only undergo thermal motion near their equilibrium position. However, a small number of molecules that are at the interface at time $t$ may enter the bulk phase at a later time. Due to the inclusion of H-bonds in the bulk phase, this scenario tends to underestimate the breaking rate of the H-bonds at the interface. This Luzar-Chander HB population-based method includes three steps as follows.
1. Sub-trajectories: Sub-trajectories with a specific length of time, $t_{\text{traj}}\sim 10$ ps are selected. Here, $t_{\text{traj}}$ is long enough to observe HB dynamics but short enough that the molecules do not complete their transition across the interface Madarasz, Rossky, and Turi (2007).
2. Sampling: At multiple, evenly spaced moments within $t_{\text{traj}}$, we identify a layer at the interface of a specified thickness $d$ (see Fig. 1) and select the water molecules within this layer for analysis. For each of these moments, the Luzar-Chandler HB population-based correlation functions Luzar and Chandler (1996a) for these molecules across the sub-trajectory $t_{\text{traj}}$ are calculated.
3. Statistics: The average properties over all sampled moments and all sub-trajectories are calculated.
For details on this LC HB operator-based method (LC method), see Appendix C.
II.2 Scenario 2: HB dynamics based on interface HB population $h^{(\text{s})}[{\mathbf{r}}(t)]$
@@
$h^{(\text{s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{j\in\mathcal{S}_{i}(t)};\\ 0&&{\text{otherwise,}}\end{array}\right.$ (4)
where ${\mathcal{S}_{i}(t)}$ denotes the set of water molecules that are H-bonded with molecule $i$ at time $t$, and $\mathcal{I}(d;t)$ is the instantaneous interface layer with thickness $d$ at time $t$, as defined previously. The definition of $h^{(\text{s})}$ combines the Luzar-Chandler’s HB population Luzar and Chandler (1993a, 1996a) $h$ and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017). Then the correlation function $c^{(\text{s})}(t)$ that describes the fluctuation of H-bonds at the interface:
$c^{(\text{s})}(t)=\langle h^{(\text{s})}(0)h^{(\text{s})}(t)\rangle/\langle h^{(\text{s})}\rangle,$ (5)
can be obtained.
Similar to functions $n(t)$ and $k(t)$ in Ref. 76 (Eq.s  12 and  13 in Appendix A), the corresponding correlation function
$n^{(\text{s})}(t)=\langle h^{(\text{s})}(0)[1-h^{(\text{s})}(t)]h^{(\text{d,s})}\rangle/\langle h^{(\text{s})}\rangle,$ (6)
@@
$k^{(\text{s})}(t)=-{dc^{(\text{s})}(t)}/{dt}$ (7)
are obtained. The $h^{(\text{d,s})}(t)$ is 1 when a tagged pair of water molecules $i$, $j$ is at the interface and the inter-oxygen distance between the two molecules is less than the cutoff radius $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise, i.e.,
@@
Therefore, $n^{(\text{s})}(t)$ represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules at the interface are unbonded but remain at the interface and separated by less than $r_{\text{OO}}^{\text{c}}$; $k^{(\text{s})}(t)$ measures the effective decay rate of H-bonds at the interface. The functions defined in Eq.s 57 are used to determine the reaction rate constants of breaking and reforming and the lifetimes of H-bonds at the interface by Luzar and Chandler (1996a); Khaliullin and Kühne (2013)
@@
In this IHB scenario, choosing the water molecules and H-bonds at the interface is accurate. However, for some special H-bonds, if it connects such two water molecules, one is at the interface and the other is in bulk water, the HB breaking reaction rate of such H-bonds will be increased. Therefore, in contrast to the LC method used in scenario 1, the IHB method used in this scenario overestimates the HB breaking rate constant.

### v3

@@ Abstract
The air-water interface plays a crucial role in many aspects of science, because of its unique properties, such as a two-dimensional hydrogen bond (HB) network and completely different HB dynamics compared to bulk water. However, accurately determining the boundary of interfacial and bulk water, that is, the thickness of the air-water interface, still challenges experimentalists. Various simulation-based methods have been developed to estimate the thickness, converging on a range of approximately 3–10 (Å). In this study, we introduce a novel approach, grounded in density functional theory-based molecular dynamics and deep potential molecular dynamics simulations, to measure the air-water interface thickness, offering a different perspective based on prior research. To capture realistic HB dynamics in the air-water interface, two extreme scenarios of the interface HB dynamics are obtained: one underestimates the interface HB dynamics, while the other overestimates it. Surprisingly, our results suggest that the interface HB dynamics in both scenarios converges as the thickness of the air-water interface increases to 4 (Å). This convergence point, indicative of the realistic interface thickness, is also validated by our calculation of anisotropic decay of OH stretch and the free OH dynamics at the air-water interface.
@@ I Introduction
Advances in the study of hydrogen bond (HB) dynamics at the air-water interface have been significant. Liu et al. Liu, Harder, and Berne (2005) used molecular dynamics (MD) simulations to demonstrate faster HB breaking and forming at the interface than bulk water, attributed to quicker translational diffusion. From sum frequency generation (SFG) vibrational spectroscopy, Gan et al. Gan et al. (2006a) found that at the air-water interface, singly hydrogen (H)-bonded water molecules align almost parallel to the interface with limited orientational variation, while doubly H-bonded donor molecules orient their dipole vectors away from the liquid phase, highlighting diverse behaviors among interfacial water molecules. Almost concurrently, through time-resolved SFG vibrational spectroscopy, McGuire and ShenMcguire and Shen (2006) observed ultrafast vibrational dynamics at the interface, noting that the relaxation behaviors of interfacially bonded OH stretch modes on subpicosecond time scales were akin to those in bulk water, encompassing spectral diffusion, vibrational relaxation, and thermalization. Pioneering work by Tahara’s group Singh et al. (2013); Inoue et al. (2016), which presented the first two-dimensional heterodyne-detected vibrational SFG (2D HD-VSFG) spectra of the OH stretch region at the interface, highlighted diverse behaviors among interfacial HB OH groups. Subsequent studies, including those by Jeon et al. Jeon et al. (2017) and Ojha and Kühne Ojha and Kühne (2021), have used MD and ab initio MD (AIMD) simulations to explore the structure and dynamics of interfacial water, uncovering weaker H-bonds and faster vibrational spectral dynamics of free OH groups compared to H-bonded OH groups at the interface. These collective insights enhance our understanding of the vibrational energy relaxation, HB dynamics, and interactions of water molecules at the air-water interface. Building upon this knowledge of interfacial behavior, significant efforts have also been directed toward quantifying the physical characteristics of the interface, including its thickness, which plays a crucial role in understanding its molecular interactions and behavior.
@@
The air-water interface thickness has been measured via ellipsometry Raman and Ramdas (1927); McBain, Bacon, and Bruce (1939); Kinosita and Yokota (1965), relative permittivity measurements Fumagalli et al. (2018), X-ray reflectivity Braslau et al. (1985, 1988), SFG spectroscopy Superfine, Huang, and Shen (1991); Du et al. (1993); Ostroverkhov, Waychunas, and Shen (2005); Groenzin et al. (2007); Shen (2012); Ishiyama, Takahashi, and Morita (2012); Shen (2013); Singh et al. (2013); Ohto et al. (2015); Tang et al. (2020), classical MD simulationsLiu, Harder, and Berne (2005); Kuo et al. (2006); Wick et al. (2007); Morita and Garrett (2008); Sedlmeier, Horinek, and Netz (2009); Fan et al. (2009); Ishiyama and Morita (2009); Willard and Chandler (2010); Vila Verde, Bolhuis, and Campen (2012); Ishiyama, Morita, and Tahara (2015); Nagata et al. (2015) and the AIMD simulationsKühne et al. (2011); Nagata et al. (2012); Kessler et al. (2015); Jeon et al. (2017); Ohto et al. (2019); Tang et al. (2020) to mention just a few. There is a consensus that the thickness of the air-water interface is about 3–10 (Å) Jungwirth and Tobias (2006); Sedlmeier, Horinek, and Netz (2009); Ishiyama and Morita (2009); Kuo et al. (2006); Wick et al. (2007); Fan et al. (2009); Kühne et al. (2011); Baer et al. (2014); Kessler et al. (2015); Nguyen, Nakahara, and Phan (2020); Peng et al. (2021). Nonetheless, accurately determining the thickness remains experimentally challenging. The MD and Monte Carlo (MC) simulations of the air-water interface yield molecular-level information not readily available in experiments. These simulations, which utilize various intermolecular potential functions, have played a crucial role in estimating the thickness Townsend, Gryko, and Rice (1985); Wilson, Pohorille, and Pratt (1987); Matsumoto and Kataoka (1988); Townsend and Rice (1991); Lie et al. (1993); Alejandre, Tildesley, and Chapela (1995); Taylor, Dang, and Garrett (1996); Dang and Chang (1997b); Tang et al. (2020). Additionally, density functional theory-based MD (DFTMD) simulations Car and Parrinello (1985); Marx and Hutter (2000); Car (2002); Kuo and Mundy (2004); Kühne et al. (2007, 2011); Khaliullin and Kühne (2013) also offer a predictive platform for understanding density profiles and determining the thickness of the air-liquid interfaces Blas et al. (2001); Kühne et al. (2011); Kessler et al. (2015); Pezzotti, Galimberti, and Gaigeot (2017). Nevertheless, DFTMD simulations are constrained by limitations in time and the number of molecules they can model. Traditional force field approaches, however, often lack the accuracy required to describe complex interface systems. Schran et al. (2021) Recently, deep potential molecular dynamics (DeePMD) simulations based on machine learning potential (MLPs) have emerged as a promising alternative, offering a solution to the accuracy-versus-efficiency dilemma in molecular simulations. Zeng et al. (2023) One of the most accurate MLPs for water is MB-pol, which accurately reproduces many properties of water across the phase diagram. Babin, Leforestier, and Paesani (2013); Babin, Medders, and Paesani (2014); Medders, Babin, and Paesani (2014); Bore and Paesani (2023); Muniz et al. (2021) Moreover, MB-pol has been used to obtain the VSFG and surface tension of air-water interface, demonstrating excellent agreement between theoretical predictions and experimental measurements. Medders and Paesani (2016)
@@
Inspired by the above experimental and simulation results, and with the motivation of capturing realistic HB dynamics at interfaces, we have designed an approach based on two extreme scenarios of interface HB dynamics by utilizing the trajectories of DFTMD and DeePMD simulations based on MB-pol. In the first scenario, for the set of molecules located in the interface layer at given sampling times, we use the Luzar-Chandler (LC) HB population operator Luzar and Chandler (1996) to obtain the HB dynamics of these interface molecules. In the second scenario, taking inspiration from Luzar and Chandler’s HB population and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017), we have developed an interface HB (IHB) population operator. This operator aims to provide a refined understanding of HB dynamics specifically at the interface.
@@
Due to the thermal motion of water molecules, two scenarios may occur. In the first scenario, the water molecules under observation might transition into the bulk phase. In the second scenario, if a water molecule resides in the interface region and its H-bonded partner moves outside the interface area, then such a pair of molecules will no longer be of concern. Based on the study of interfacial water dynamics by Liu et al. Liu, Harder, and Berne (2005), Gan et al. Gan et al. (2006a), Singh et al. Singh et al. (2013) and Jeon et al. Jeon et al. (2017), as well as the investigation into the time-dependent spectral evolution of H-bonded and free water molecules by Ojha and Kühne Ojha and Kühne (2021), our approach indicates that the HB dynamics derived from the first scenario will be slower the genuine interfacial HB dynamics. Conversely, the one obtained from the second scenario will exhibit faster dynamics than the genuine one.
Building upon the foundation laid by methods reliant on the density criterion Townsend and Rice (1991); Beaglehole and Wilson (1993); Lie et al. (1993); Taylor, Dang, and Garrett (1996); Sokhan and Tildesley (1997); Alejandre, Tildesley, and Chapela (1995); Morita and Hynes (2000); Paul and Chandra (2004); Morita (2006); Zhao et al. (2018), our approach introduces an alternative way of determining interface thickness through the analysis of the convergence of interfacial HB dynamics properties. This approach effectively bypasses the necessity of accounting for liquid density. As such, it offers another perspective for measuring the thickness of the air-water interface. Furthermore, the principles underlying our approach hold potential for application to a broader range of systems, such as solution interfaces and ion shells, offering a flexible tool for interface studies.
@@ II Methods
Figure 1: Slab of water containing 512 water molecules with the instantaneous surface $\mathcal{S}$ represented as a blue mesh on the upper and lower phase boundary. Grey surface, which represents an imaginary interface $\mathcal{S}^{\prime}$, is obtained by translating the surface to the inside of the system along the normal ($z$-axis) by distance $d$. This variable $d$ is then utilized to measure the thickness of the air-water interface.
For the slab in the cuboid simulation box, an imaginary surface $\mathcal{S}^{\prime}(t)$ is obtained by translating the surface $\mathcal{S}(t)$ along the system’s normal (into bulk) to a distance $d$. The region between the two surfaces $\mathcal{S}(t)$ and $\mathcal{S}^{\prime}(t)$ is defined as the air-water interface. Below we will combine two extreme scenarios to investigate the HB dynamics at the instantaneous air-water interface.
II.1 Scenario 1: HB Dynamics Based on the Luzar-Chandler HB Population
As the first scenario we use the Luzar-Chandler HB population and employ a technique that samples water molecules right at the instantaneous interface for certain sampling time points. In Scenario 1, we divide the simulation trajectory into multiple subtrajectories of length $t_{\rm traj}$. Within these subtrajectories, the majority of water molecules exhibit thermal fluctuations near their equilibrium positions. Meanwhile, a select group of molecules initially at the interface may transition into the bulk phase at a later time. Due to the inclusion of H-bonds in the bulk phase, this scenario tends to underestimate the breaking rate of the H-bonds at the interface. In Section 3 of this paper, we will see this result, combined with the outcome from Scenario 2, can be used to estimate the thickness of the air-water interface. This method includes three steps as follows:
(a) Sub-trajectories. Sub-trajectories with a specific length of time, $t_{\text{traj}}$, are selected. In this work, $t_{\rm traj}$ is set to be $40$ (ps), which is long enough to observe HB dynamics but short enough that not all the molecules complete their transition across the interface Madarász, Rossky, and Turi (2007).
(b) Sampling. For each time step, we identify a pair of the air-water interfaces of a specified thickness $d$ as shown in Figure 1. At evenly spaced moments within $t_{\text{traj}}$, we select the water molecules within the interfaces. For the union of the interfacial water molecules picked at all these time moments, the Luzar-Chandler HB population-based correlation functions Luzar and Chandler (1996) across the sub-trajectory are calculated.
(c) Statistics. The average correlation functions across all sub-trajectories are calculated.
In this scenario, the computational procedure for calculating HB dynamics follows the same methodology as that used in works based on the Luzar-Chandler’s method (LC method)Luzar (2000); Benjamin (2005); Liu, Harder, and Berne (2005). For details on this method, please refer to the Supporting Information.
II.2 Scenario 2: HB Dynamics Based on Interface HB Population
@@
$h^{(\text{s})}[{\mathbf{r}}(t)]=\left\{\begin{array}[]{rcl}1&&{{\mathbf{r}}_{i}\in\mathcal{I}(d;t),{\mathbf{r}}_{j}\in\mathcal{I}(d;t)},\\ &&{j\in\mathcal{B}_{i}(t)};\\ 0&&{\text{otherwise}}\end{array}\right.$ (4)
where ${\mathbf{r}}(t)$ is the configuration of the system at time $t$, ${\mathbf{r}}_{i}$ is the position coordinate of the oxygen atom in the $i$-th water molecule, ${\mathcal{B}_{i}(t)}$ denotes the set of water molecules that are H-bonded with molecule $i$ at time $t$, and $\mathcal{I}(d;t)$ is the instantaneous interface layer with thickness $d$ at time $t$. The definition of $h^{(\text{s})}$ combines the Luzar-Chandler’s HB population Luzar and Chandler (1993, 1996) $h$ and the characteristic function introduced by Giberti and Hassanali Giberti and Hassanali (2017). Then the correlation function $c^{(\text{s})}(t)$ that describes the fluctuation of H-bonds at the interface:
$c^{(\text{s})}(t)=\frac{\langle h^{(\text{s})}(0)h^{(\text{s})}(t)\rangle}{\langle h^{(\text{s})}\rangle},$ (5)
can be obtained. Similar to functions $n(t)$ and $k(t)$ in ref 84 (eqs 1 and 2 in Supporting Information), the corresponding correlation function
$n^{(\text{s})}(t)=\frac{\langle h^{(\text{s})}(0)[1-h^{(\text{s})}(t)]h^{(\text{d,s})}\rangle}{\langle h^{(\text{s})}\rangle},$ (6)
@@
$k^{(\text{s})}(t)=-\frac{\mathrm{d}c^{(\text{s})}(t)}{\mathrm{d}t}$ (7)
are obtained. The $h^{(\text{d,s})}(t)$ is 1 when a tagged pair of water molecules $i$, $j$ is at the interface and the interoxygen distance between the two molecules is less than the cutoff radius $r^{\mathrm{c}}_{\mathrm{OO}}$ at time $t$, and 0 otherwise, i.e.,
@@
Therefore, $n^{(\text{s})}(t)$ represents the probability at time $t$ that a tagged pair of initially H-bonded water molecules at the interface are unbonded but remain at the interface and separated by less than $r_{\text{OO}}^{\text{c}}$; $k^{(\text{s})}(t)$ measures the effective decay rate of H-bonds at the interface. The functions defined in eqs 57 are used to determine the reaction rate constants of breaking and reforming and the lifetimes of H-bonds at the interface by Luzar and Chandler (1996); Khaliullin and Kühne (2013)
@@
In this IHB scenario, choosing the water molecules and H-bonds at the interface is accurate. However, for some special H-bonds, if it connects such two water molecules, one is at the interface and the other is in the bulk phase, the HB breaking reaction rate of such H-bonds will be increased. Therefore, in contrast to the LC method used in Scenario 1, the IHB method used in this scenario overestimates the HB breaking rate constant.